OpenStudy (rock_mit182):

Solve the compound inequalitie: 0 ≤ 4 + 2x ≤ 22

OpenStudy (rock_mit182):

i have to substract -4 from both sides, right ?

OpenStudy (zarkon):

you subtract 4 from all 3 pieces

OpenStudy (anonymous):

yes, from all the sides

OpenStudy (rock_mit182):

and then divide by two in all sides ?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

correct

OpenStudy (rock_mit182):

\[-4+0\le -4+4+2x \le22-4\]

OpenStudy (anonymous):

yes then combine like terms

OpenStudy (rock_mit182):

\[-4\le2x \le18\]

OpenStudy (anonymous):

yup, then divide by 2

OpenStudy (rock_mit182):

\[\frac{ -4 }{ 2 }\le \frac{ 2x }{ 2}\le \frac{ 18 }{ 2 }\]

OpenStudy (rock_mit182):

\[-2\le x \le 9\]

OpenStudy (rock_mit182):

when the signs switch?

OpenStudy (anonymous):

the sings switch only when dividing or multiplying by a negative number

OpenStudy (anonymous):

because the 2 is not negative, the sings don't switch

OpenStudy (anonymous):

so that is the right answer, good job

OpenStudy (rock_mit182):



OpenStudy (rock_mit182):

is this the interval notation ?

OpenStudy (anonymous):

correct

OpenStudy (rock_mit182):

what kind of solution is this ? union, intersection, and, or,

OpenStudy (rock_mit182):

i guess is a compound inequality...

OpenStudy (rock_mit182):

help me out please !

OpenStudy (anonymous):

im not sure on that part sorry